Calculate the rest energy of a single proton. Use the equation E=mc² to solve the problem.

1
Expert's answer
2022-05-30T15:18:19-0400
"E=mc^2\\
=\rm 1.67*10^{-27}\: kg\times (3*10^8\:\rm m/s)^2=1.5*10^{-10}\: J"